(√(1-y^2)dx)-(√(1-x^2)dy)=0, y(0)=√3/2)
1
Expert's answer
2020-12-02T20:07:31-0500

"(\sqrt{1-y^2})dx-(\sqrt{1-x^2})dy=0\\
(\sqrt{1-y^2})dx=(\sqrt{1-x^2})dy\\
\frac{1}{\sqrt{1-y^2}}dy=\frac{1}{\sqrt{1-x^2}}dx\\
\text{Integrate both sides}\\
\sin^{-1}(y)=\sin^{-1}(x)+c\\
\text{But} y(0)=\frac{\sqrt{3}}{2}\\
\sin^{-1}(\frac{\sqrt{3}}{2})=\sin^{-1}(0)+c\\
60^{\circ}=0+c\\
c=60^{\circ}\\
\text{Therefore}, \\
\sin^{-1}(y)=\sin^{-1}(x)+60°"